OpenStudy (luigi0210):

Optimization:

OpenStudy (luigi0210):

If \(1,200 cm^2\) of material is available to make a box with a square base and an open top, find the largest possible volume of the box.

OpenStudy (dumbcow):

Let x be length and width, h be height SA = x^2 + 4hx = 1200 --> h = (1200 -x^2)/4x V = hx^2 = 300x - x^3/4 maximize volume by setting derivative equal to 0 dV/dx = 300 - 3x^2/4 = 0 --> x = 20 --> max V = 4000

OpenStudy (luigi0210):

Thank you :)

OpenStudy (dumbcow):

yw :)